0 A consumers budget set for two goods (Xand is 500 z5X+10Y. a. The budget set is illustrated below. What are the values of A and B? Good Y Book Good X < Prev 20 15. Next> MacBook Air
b. Does the budget set change if the prices of both goods double and the consumers income also d O Yes, it shifts in toward the origin. O No, it does not change. O Yes, it shifts out from the origin. O Yes, it rotates clockwise. c. Given the equation for the budget set, what are the prices of the two goods? What is the consumers income? < Prev 2 of 15 Next> MacBook Air 漪: od FA O f3 F2
Answer #1

a)

A represents the intercept of good Y and B represents the intercept of good X.

Thus for finding A, we will keep the value of X in the budget equation equal to zero and since A is intercept of Y it will be A in the equation in place of Y. Budget equation will become -

500 = 5*0 + 10 (A)

500 / 10 = A

A = 50

For finding B, we will keep the value of Y in the budget equation equal to zero and since B is intercept of X it will be B in the equation in place of X.

500 = 5 (B) + 10*0

500 / 5 = B

B = 100

b)

Answer - No, it does not change

Reason - When the prices of both goods double consumer will be able to purchase less of each good but with doubling of the income it will come back to the original budget line purchasing the previous combination and budget lines stays the same.

c)

Budget line - 500 = 5X + 10Y

The general equation of the budget line is - M = Px (X) + Py (Y)

Here M represents income of the consumer. Px represents the price of good X and Py represents price of good Y.

Thus given the budget line -

500 - Consumer's income

5 - Price of good X

10 - Price of good Y